OpenStudy (anonymous):

Calculate the line integral where; F(x, y, z) = (yz, xz, xy) Υ(t)= (2 cos t, t^2 sin t, cos^2 t) tε [0; 3] HINT: it is a conservative function

OpenStudy (anonymous):

I believe they want you to solve some exact equations for this one, if they give you the hint that it's conservative, that means that the line integral will only depend on it's initial and end point. So that will means that: \[\int\limits_C yzdx+xzdy+xydz=\int\limits_a^b \nabla fdr\]

OpenStudy (anonymous):

Do you know how to setup your equations from here?

OpenStudy (anonymous):

I got f (x,y,z) = xyz , not sure about doing the integration....

OpenStudy (anonymous):

So you have found your potential function correctly, now you need to evaluate that function from the beginning point to the end point. You can also rewrite your potential function by substituting your given vector function.

OpenStudy (anonymous):

\[F(x,y,z)=xyz=2\cos (t) t^2 \sin(t) \cos^2t =2t^2 \cos^3(t)\sin(t)\]

OpenStudy (anonymous):

from t=0 to t=3

OpenStudy (anonymous):

I was told when the function is conservative it is a path independant.......but looks like for the final answer we still should Integrate the path by pluging the xyz value and using the point. right??

OpenStudy (anonymous):

No if you found your potential function then you just need to evalaute that function from the end and the start point. So if you have found your potential function, then there's is nothing much more to do than that.

OpenStudy (anonymous):

in this case, please let me know what you got for the final answer?

OpenStudy (anonymous):

18cos^3(3)sin(3)